OpenStudy (anonymous):

Solve D = 3(2h – k) for h

Parth (parthkohli):

Start by dividing both sides by 3.

OpenStudy (anonymous):

D = 3(2h - k) 6h - k = D 6h = D + k h = (D + k)/6 Answer: h = (D + k)/6 OR h = 1/6D + 1/6k

Parth (parthkohli):

Then, add k to both sides and divide both sides by 2.

Parth (parthkohli):

\(\Rightarrow \Large {D \over 3}= 2h - k \) @sambruce0825 Do you get the concept till here?

OpenStudy (anonymous):

Not exactly.

Parth (parthkohli):

Then, try to add k to both sides. \(\Rightarrow \Large {D \over 3}+ k = 2h \)

OpenStudy (lgbasallote):

@Master.RohanChakraborty you failed to distribute 3 into k ;)

Parth (parthkohli):

You have to isolate a variable, step by step.

OpenStudy (anonymous):

will you do it @lgbasallote

Parth (parthkohli):

\(\Rightarrow h = \Huge{ {D \over 3} + k\over 2} \)

OpenStudy (lgbasallote):

kohli is already doing it lol. it's not my habit to answer a question that someone is already entertaining <tips hat>

Parth (parthkohli):

\(\Rightarrow \Large h = {D \over 6 } + {k \over 2}\) Step by step division.

OpenStudy (anonymous):

okay Leave that aside guys @ParthKohli @lgbasallote Do u understand the question @sambruce0825 ??? Concept is @ParthKohli

Parth (parthkohli):

4 tags at once... impressive #SHAKA#

OpenStudy (anonymous):

Would that be the same as D/3 + k?

Parth (parthkohli):

No, k/2 actually

OpenStudy (anonymous):

So far none of these answers are choices.

OpenStudy (anonymous):

a medal plzzz lol J.Joking

OpenStudy (anonymous):

for the shaka

Parth (parthkohli):

Why are you asking a medal for me?

OpenStudy (anonymous):

shake it forget it